OpenStudy (anonymous):

Simplify the following expression. 3x - 6y + 2z - 5x - 3y + 8z

OpenStudy (anonymous):

3x - 6y + 2z - 5x - 3y + 8z = 3x - 5x - 6y - 3y + 2z + 8z = - 2x - 9y + 10z

OpenStudy (anonymous):

basically we have to take like terms and resolve them....

OpenStudy (turingtest):

combine like terms (i.e. add the x's y's and z's together): =3x-5x-6y-3y+2z+8z =-2x-9y+10z sorry harkirat I was almost done though. Don't mean to echo you

OpenStudy (anonymous):

no issues, many a times i keep clicking the POST button but my post goes thru much later....